# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: I. Inherited and genuine singularities

### v1

@@ IV.1.2 Linear configuration with different damping and/or amplification rates for neighbor modes
$\gamma_{+}$$\pm\alpha_{+}$ $\langle b_{1}\rangle$, $\langle b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1x2 2x2 1x2 2x2
$\langle b_{2}\rangle$, $\langle b_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1 1x2 1x2
$\alpha_{-}$ $\langle b_{3}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle b_{4}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1
$-\alpha_{-}$ $\langle b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\rangle$ 1
$2\gamma_{+}$$\pm 2\alpha_{+}$ $\langle b_{1}^{2}\rangle$, $\langle b_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 4x4 1x3 2x3
$\alpha_{+}-\alpha_{+}$ $\langle b_{1}^{\dagger}b_{1}\rangle$ 2 +
±$\pm 2\alpha_{+}$ $\langle b_{2}^{2}\rangle$, $\langle b_{2}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{+}-\alpha_{+}$ $\langle b_{2}^{\dagger}b_{2}\rangle$ 2
±$\pm\alpha_{+}$ $\langle b_{1}b_{2}\rangle$, $\langle b_{1}^{\dagger}b_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{2}}\rangle$ 2 2x4 1x4 1x4
$\alpha_{+}-\alpha_{+}$ $\langle b_{1}^{\dagger}b_{2}\rangle$, $\langle b_{1}b_{2}^{\dagger}\rangle$ 2
$\alpha_{-}\pm\alpha_{+}$ $\langle b_{1}b_{3}\rangle$, $\langle b_{1}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{3}}\rangle$ 2 2x2 16x2 2x2 16x2
$\langle b_{1}b_{4}\rangle$, $\langle b_{1}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{4}}\rangle$ 2 2x2 2x2
$\alpha_{-}\pm\alpha_{+}$ $\langle b_{2}b_{3}\rangle$, $\langle b_{2}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{3}}\rangle$ 2 2x2 2x2
$\langle b_{2}b_{4}\rangle$, $\langle b_{2}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{4}}\rangle$ 2 2x2 2x2
$-\alpha_{-}\pm\alpha_{+}$ $\langle b_{1}b_{3}^{\dagger}\rangle$, $\langle b_{1}^{\dagger}b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{5}}\rangle$ 2 2x2 2x2
$\langle b_{1}b_{4}^{\dagger}\rangle$, $\langle b_{1}^{\dagger}b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{6}}\rangle$ 2 2x2 2x2
$-\alpha_{-}\pm\alpha_{+}$ $\langle b_{2}b_{3}^{\dagger}\rangle$, $\langle b_{2}^{\dagger}b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{5}}\rangle$ 2 2x2 2x2
$\langle b_{2}b_{4}^{\dagger}\rangle$, $\langle b_{2}^{\dagger}b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{6}}\rangle$ 2 2x2 2x2
$2\alpha_{-}$ $\langle b_{3}^{2}\rangle$, $\langle b_{4}^{2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$, $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle b_{3}b_{4}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{4}}\rangle$ 2
$-2\alpha_{-}$ $\langle b_{3}^{\dagger 2}\rangle$, $\langle b_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$, $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle b_{3}^{\dagger}b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{6}}\rangle$ 2
$\alpha_{-}-\alpha_{-}$ $\langle b_{3}^{\dagger}b_{3}\rangle$, $\langle b_{4}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{3}}\rangle$, $\langle\bm{\hat{B}_{6}\hat{B}_{4}}\rangle$ 2 8x1 8x1 4x1 4x1
$\langle b_{3}^{\dagger}b_{4}\rangle$, $\langle b_{3}b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{4}}\rangle$, $\langle\bm{\hat{B}_{3}\hat{B}_{6}}\rangle$ 2
@@
Table 4: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(4)}_{\rm l2}}$, given in Eq. (44) with Eq. (49) for the linear four-mode bosonic system with different damping and/or amplification rates of neighbor modes, as derived from the equations for the FOMs up to second order assuming $\alpha_{+}=0$. For details, see the caption to Tab. 1.
IV.2 Circular configuration
[…]
Whereas the Heisenberg-Langevin equations keep the form of Eq. (39), the original dynamical matrix $\bm{M^{(4)}_{\rm l}}$ of the linear configuration in Eq. (44) is modified into the form
We derive the eigenvalues of the matrix $\bm{M^{(4)}_{\rm c}}$ under the condition of the equal damping and/or amplification rates of the neighbor modes valid for the configuration plotted in Fig. 1(d); i.e.,
We obtain the following eigenvalues
$\lambda_{1,2}^{M^{(4)}_{\rm c1}}$ $=$ $-i\gamma_{+}-\alpha_{\pm},$
$\lambda_{3,4}^{M^{(4)}_{\rm c1}}$ $=$ $-i\gamma_{+}+\alpha_{\mp}$ (60)
that are accompanied by the following eigenvectors
For $\beta=0$, we have $\alpha_{+}=-\alpha_{-}$ and $\chi^{*}=-\chi$. This leads to the relations $\lambda_{1}^{M^{(4)}_{\rm c1}}=\lambda_{3}^{M^{(4)}_{\rm c1}}$ and $\lambda_{2}^{M^{(4)}_{\rm c1}}=\lambda_{4}^{M^{(4)}_{\rm c1}}$. Also the corresponding eigenvectors coincide: $\bm{y_{1}^{M^{(4)}_{\rm c1}}}=\bm{y_{3}^{M^{(4)}_{\rm c1}}}$ and $\bm{y_{2}^{M^{(4)}_{\rm c1}}}=\bm{y_{4}^{M^{(4)}_{\rm c1}}}$. As a consequence, the two QEPs with second-order EDs occur. This means that the two QHPs with second-order EDs and DDs for the $8\times 8$ matrix $\bm{M^{(4)}_{\rm c1}}$ are formed. These QHPs occur under the condition specified in Eq. (22) that identifies the QHPs in the analyzed two-mode bosonic system. The real values of the eigenvalues $\lambda_{j}^{M^{(4)}_{\rm c1}}$ for $j=1,\ldots,4$ are plotted in Fig. 2(e) together with the condition for the QHPs.
Using the eigenvalues and eigenvectors of the $4\times 4$ matrix $\bm{M^{(4)}_{\rm c1}}$ and the $2\times 2$ matrix $\bm{\xi}$ given in Eqs. (60), (61), (13), and (14), we arrive at the system dynamical matrix in its diagonal form (for details, see Appendix A). This brings us the new field operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{2},\hat{b}_{3}^{\dagger},\hat{b}_{4}^{\dagger},\hat{b}_{3},\hat{b}_{4},\hat{b}_{1}^{\dagger},\hat{b}_{2}^{\dagger}\right]^{T}$ suitable for revealing QEPs and QHPs found in the dynamics of FOMs. As the structure of eigenvalues and eigenvectors is the same as that characterizing the four-mode linear system with equal damping and/or amplification rates of neighbor modes, the appropriate QEPs and QHPs are given in Tab. 3.
V Five-mode bosonic systems
V.1 Linear configuration
[…]
$\frac{d\hat{\bm{a}}}{dt}$ $=$ $-i\bm{M^{(5)}_{\rm l}}\hat{\bm{a}}+\hat{\bm{L}}.$ (63)
The dynamical matrix $\bm{M^{(5)}_{\rm l}}$ introduced in Eq. (63) is written with the help of $2\times 2$ submatrices introduced in Eq. (10) as
Motivated by $\mathcal{PT}$ symmetry we assume:
$2\gamma_{3}=\gamma_{12}+\gamma_{45}.$ (71)
Then, diagonalization of the matrix $\bm{M^{(5)}_{\rm l}}$ results in the following eigenvalues ${\lambda_{j}^{M^{(5)}_{\rm l}}}$:
$\lambda_{1}^{M^{(5)}_{\rm l}}$ $=$ $-i\gamma_{+},$
$\lambda_{2,3}^{M^{(5)}_{\rm l}}$ $=$ $-i\gamma_{+}\pm\alpha_{-},$
$\lambda_{4,5}^{M^{(5)}_{\rm l}}$ $=$ $-i\gamma_{+}\pm\alpha_{+}.$ (72)
The corresponding eigenvectors are derived as follows:
$\frac{\kappa^{2}}{\epsilon^{2}}+\frac{4\gamma_{-}^{2}}{\epsilon^{2}}=1.$ (74)
The real parts of the eigenvalues $\lambda_{j}^{M^{(5)}_{\rm l}}$ forming QHPs ($j=2,\ldots,5$) are plotted in Fig. 2(f).
The eigenvalues and eigenvectors of the $10\times 10$ matrix $\bm{M^{(5)}_{\rm l}}$ constructed from the formulas given in Eqs. (72) and (73) together with those in Eqs. (13) and (14) (for details, see Appendix A) allow us to describe the system dynamics via a diagonal dynamical matrix. The appropriate operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{3},\hat{b}_{2}^{\dagger},\hat{b}_{3}^{\dagger},\hat{b}_{4},\hat{b}_{5},\hat{b}_{4}^{\dagger},\hat{b}_{5}^{\dagger}\right]^{T}$ then reveal QEPs and QHPs found in the dynamics of the first- and second-order FOMs (see Tab. 5).

### v3

@@ 0 $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x9 1x6
$\gamma_{+}$ $\pm\alpha_{+}$ $\langle\hat{b}_{1}\rangle$, $\langle\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1x2 2x2 1x2 2x2
$\alpha_{-}$ $\langle\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1
$\alpha_{-}$ $\langle\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\rangle$ 1
$2\gamma_{+}$ $\pm 2\alpha_{+}$ $\langle\hat{b}_{1}^{2}\rangle$, $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 4x4 1x3 2x3
$\alpha_{+}-\alpha_{+}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{1}\rangle$ 2 +
$\pm 2\alpha_{+}$ $\langle\hat{b}_{2}^{2}\rangle$, $\langle\hat{b}_{2}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{+}-\alpha_{+}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{2}\rangle$ 2
$\pm\alpha_{+}$ $\langle\hat{b}_{1}\hat{b}_{2}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{2}}\rangle$ 2 2x4 1x4 1x4
$\alpha_{+}-\alpha_{+}$ $\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{1}\hat{b}_{2}^{\dagger}\rangle$ 2
$\alpha_{-}\pm\alpha_{+}$ $\langle\hat{b}_{1}\hat{b}_{3}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{3}}\rangle$ 2 2x2 16x2 2x2 16x2
$\langle\hat{b}_{1}\hat{b}_{4}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{4}}\rangle$ 2 2x2 2x2
$\alpha_{-}\pm\alpha_{+}$ $\langle\hat{b}_{2}\hat{b}_{3}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{3}}\rangle$ 2 2x2 2x2
$\langle\hat{b}_{2}\hat{b}_{4}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{4}}\rangle$ 2 2x2 2x2
$\alpha_{-}\pm\alpha_{+}$ $\langle\hat{b}_{1}\hat{b}_{3}^{\dagger}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{5}}\rangle$ 2 2x2 2x2
$\langle\hat{b}_{1}\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{1}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{6}}\rangle$ 2 2x2 2x2
$\alpha_{-}\pm\alpha_{+}$ $\langle\hat{b}_{2}\hat{b}_{3}^{\dagger}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{5}}\rangle$ 2 2x2 2x2
$\langle\hat{b}_{2}\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{6}}\rangle$ 2 2x2 2x2
$2\alpha_{-}$ $\langle\hat{b}_{3}^{2}\rangle$, $\langle\hat{b}_{4}^{2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$, $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle\hat{b}_{3}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{4}}\rangle$ 2
$2\alpha_{-}$ $\langle\hat{b}_{3}^{\dagger 2}\rangle$, $\langle\hat{b}_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$, $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{6}}\rangle$ 2
$\alpha_{-}-\alpha_{-}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{3}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{3}}\rangle$, $\langle\bm{\hat{B}_{6}\hat{B}_{4}}\rangle$ 2 8x1 8x1 4x1 4x1
$\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{3}\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{4}}\rangle$, $\langle\bm{\hat{B}_{3}\hat{B}_{6}}\rangle$ 2
@@
Table 4: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(4)}_{\rm l2}}$, given in Eq. (49) with Eq. (54) for the linear four-mode bosonic system with different damping and/or amplification rates of neighbor modes, as derived from the equations for the FOMs up to second order assuming $\alpha_{+}=0$. We have $\bm{\hat{B}_{j}}\equiv[\hat{b}_{j},\hat{b}_{j}^{\dagger}]$ for $j=1,2$, $\bm{\hat{B}_{j}}\equiv[\hat{b}_{j}]$ for $j=3,4$, $\bm{\hat{B}_{5}}=\bm{\hat{B}_{3}}^{\dagger}$, $\bm{\hat{B}_{6}}=\bm{\hat{B}_{4}}^{\dagger}$ and more details are given in the caption to Tab. 1.
In the linear five-mode bosonic system described by the dynamical matrix ${\bm{M}}_{\rm l}^{(5)}$ given in Eq. (74), with damping and/or amplification rates satisfying the conditions in Eqs. (75) and (76), the QEPs and QHPs observed in the dynamics of first- and second-order FOMs are summarized in Tab. 5. Notably, two inherited QEPs with second-order EDs give rise to genuine QEPs exhibiting second-, third-, and fourth-order EDs in various configurations within the spectral structure of second-order FOMs.